If in a triangle ,tanA.tanB=2,
Then find the value of cosA.cosB/cosC

Grade:11

1 Answers

Arun
6 years ago
you can write-
sinA sinB= 2 cosA cosB
now sinA sinB –cosA cosB= cosA cosB
thus cos (A+B)= cosA cosB
now in a triangle, A+B= \pi–C
thus cos(\pi–C) = cosA cosB
so cosA*cosB/cosC= – 1